OpenStudy (anonymous):

I need to convert this to basic SI form: -0.000398*dm^3*h*km^-2. Wolfram alpha gives -1.433×10^-9 m*s but i keep gettin -10^-9. where do i go wrong? -0.000398*m^3*10^-3*3600*s*m^-2*10^-6

OpenStudy (anonymous):

I'm going to separate the numbers from the units and from the powers of ten first. (-0.000398)*(3600)*(m^3)*(s)*(m^-2)*(10^-3)*(10^-6)

OpenStudy (anonymous):

Multiplying the numbers gives -1.4328 Multiplying the units gives m*s Multiplying the powers of ten gives 10^-9

OpenStudy (anonymous):

Oh, hold up, I missed the original expression . . . Let me start over. :">

OpenStudy (anonymous):

i did get 1.4328 too at some point but didn't know how to continue from there

OpenStudy (anonymous):

-0.000398 = -3.98×10^-4 dm^3 = (m ×10^-1)^3 = m ×10^-3 h = 3.6×10^3 s km^-2 = (m ×10^3)^-2 = m ×10^-6

OpenStudy (anonymous):

Sorry, should be m^3 ×10-3 for the dm^3 and m^-2 ×10^-6 for the km^-2.

OpenStudy (anonymous):

OK = (-3.98)*(3.6)*(m^3)*(m^-2)*(s)*(10^-4)*(10^-3)*(10^3)*(10^-6)

OpenStudy (anonymous):

= -14.328 ×10^-10 m*s = -1.4328 ×10^-9 m*s.

OpenStudy (anonymous):

thanks! now i get it.